# Does Object Binding Naturally Emerge in Large Pretrained Vision Transformers?

### v1

@@ Abstract
Object binding, the brain’s ability to bind the many features that collectively represent an object into a coherent whole, is central to human cognition. It groups low-level perceptual features into high‑level object representations, stores those objects efficiently and compositionally in memory, and supports human reasoning about individual object instances. While prior work often imposes object-centric attention (e.g., Slot Attention) explicitly to probe these benefits, it remains unclear whether this ability naturally emerges in pre-trained Vision Transformers (ViTs). Intuitively, they could: recognizing which patches belong to the same object should be useful for downstream prediction and thus guide attention. Motivated by the quadratic nature of self-attention, we hypothesize that ViTs represent whether two patches belong to the same object, a property we term IsSameObject. We decode IsSameObject from patch embeddings across ViT layers using a similarity probe, which reaches over 90% accuracy. Crucially, this object-binding capability emerges reliably in self-supervised ViTs (DINO, MAE, CLIP), but markedly weaker in ImageNet-supervised models, suggesting that binding is not a trivial architectural artifact, but an ability acquired through specific pretraining objectives. We further discover that IsSameObject is encoded in a low-dimensional subspace on top of object features, and that this signal actively guides attention. Ablating IsSameObject from model activations degrades downstream performance and works against the learning objective, implying that emergent object binding naturally serves the pretraining objective. Our findings challenge the view that ViTs lack object binding and highlight how symbolic knowledge of “which parts belong together” emerges naturally in a connectionist system.  [^1] Code available at: https://github.com/liyihao0302/vit-object-binding.
@@ 1 Introduction
Object binding has received little attention in mainstream AI research. Cognition‑inspired models [6, 7] build in human-like object‑based attention. By contrast, mainstream vision models are assumed to implicitly learn to handle multiple objects from the training set, yet empirical studies show they often "attend" only to the most salient regions and overlook the rest [8]. ViTs compute attention scores from feature‑level similarities, which in principle should help with object binding. But experiments show that self-attention behaves more like a similarity-based grouping mechanism than selective attention, which groups image patches by low-level feature similarity (e.g. color or texture), rather than isolating distinct objects [9]. Object‑centric methods like Slot Attention [10] fix this by allocating a small set of learnable slots that compete for token features, enforcing binding by design. However, whether AI vision models, especially leading ViTs, can achieve robust object binding without explicit mechanisms remains an open question.
Cognitive scientists have questioned whether ViTs can bind objects at all: arguing that they lack mechanisms for dynamically and flexibly grouping features [5]; they lack recurrence necessary for iterative refinement of object representations in the cognitive science field [11, 7]; and as purely connectionist models, they appear incapable of true symbolic processing [8]. However, these architectural limitations do not preclude binding from emerging through learning. If a model encodes whether two patches belong to the same object (IsSameObject), this signal can guide attention and improve prediction [12, 13]. Human-labeled data also reflects object-level structure, so ViTs can acquire binding by imitation. This suggests that ViTs may learn to bind objects directly from large-scale training data, without requiring explicit architectural inductive biases.
Here, we ask whether object binding naturally emerges in large, pretrained Vision Transformers, which is a question that matters for both cognitive science and AI. We propose IsSameObject (whether two patches belong to the same object) and show that it is reliably decodable (with 90.20% accuracy) using a quadratic similarity probe starting from mid-layers of the transformer layers. This effect is robust across self-supervised models such as DINO, MAE, and CLIP, but largely absent in ImageNet-supervised ViTs, suggesting that binding is an acquired ability rather than a trivial architectural artifact. Across the ViT’s layer hierarchy, it progressively encodes IsSameObject in a low-dimensional projection-space on top of the features of the object, and it guides self-attention. Ablating IsSameObject from model activations hurts downstream performance and works against the pretraining objective.
@@ 2 Related Work
Motivated by how humans naturally reason about individual objects, Object-Centric Learning (OCL [10]) aims to represent a scene as a composition of disentangled object representations. While segmentation only partitions an image into object masks, OCL goes further by encoding each object into its own representation [33]. Unsupervised approaches such as MONet [34], IODINE [33], and especially Slot Attention [10] encode scenes into a small, permutation‑invariant set of “slots” that are iteratively refined, producing robust object representations on both synthetic [10, 35] and real‑world data [36, 37] and enabling compositional generation and manipulation [38, 39, 40]. However, slot‑based models impose a fixed slot budget and require multiple refinement iterations that slows inference, and because Slot Attention is bolted onto rather than built into the transformer, scaling and training become harder [41]. Other explicit object-centric approaches include Tensor Product Representations [42] and Capsule Networks [43].
Instead of object-centric approaches that explicitly enforce object-level attention, we propose an alternative view that ViTs may already encode implicit object‑level structure. Prior work has assumed this and attempted to group patches into objects directly from activations or attention maps ViTs, using methods like clustering [44] or GraphCut  [45]. Other studies design self‑supervised objectives (e.g., object “movability” [46] or patch‑level contrastive learning [47]) and train models to strengthen object‑level grouping. In contrast, our study directly validates this assumption by showing that ViT patch embeddings intrinsically encode whether any two patches belong to the same object.
@@ 3.1 Probing IsSameObject representations
Transformers propagate information across layers according to the above equations 12, where $h^{(\ell)}$ is the residual‑stream output (which we’ll call the patch embedding) at layer $\ell$.
@@
$\textit{IsSameObject}\bigl(x_{i}^{(\ell)},x_{j}^{(\ell)}\bigr)=\phi\bigl(x_{i}^{(\ell)},x_{j}^{(\ell)}\bigr),\quad\phi:\mathbb{R}^{d}\times\mathbb{R}^{d}\to[0,1],$
@@
We consider the following hypotheses about how IsSameObject may be encoded in the model’s activations:
@@
  • It is a pairwise relationship versus a pointwise mapping (i.e. the model first maps each patch to an object identity or class, then compares).
  • The model tells objects apart using only broad class labels or object identities–i.e., it may solve binding by recognizing “this is a dog vs. this is a chair” without actually representing which specific pixels belong to each object. Such a solution is possible because class labels already provide a coarse notion of object identity without requiring pixel-level grounding.
@@
To test these hypotheses, we decode $\mathrm{IsSameObject}$ using several probe architectures, each defined by a learnable parameter matrix $W$:
1. Linear probe: $\phi_{\mathrm{lin}}(x,y)\;=\;Wx\;+\;Wy$, where $W\in\mathbb{R}^{1\times d}$.
2. Diagonal quadratic probe (specialized dims) $\phi_{\mathrm{diag}}(x,y)\;=\;x^{\top}\operatorname{diag}(w)\,y$, with $w\in\mathbb{R}^{d}$.
3. Quadratic probe (distributed) $\phi_{\mathrm{quad}}(x,y)=\;x^{\top}W^{\top}W\,y,\quad W\in\mathbb{R}^{k\times d},\;k\ll d.$
4. Object‑class/identity probes (pointwise) We first map each embedding $x,y$ to class distributions $p=\mathrm{softmax}(W_{c}x)$, $q=\mathrm{softmax}(W_{c}y)$ and train $W_{c}$ with multiclass cross‑entropy on object‑class labels (and similarly $W_{N}$ on object‑identity labels). The pointwise IsSameObject score is then
$\phi_{\mathrm{class/identity}}(x,y)=p^{\top}q=\sum_{c=1}^{N_{c}}p(c)\,q(c)=\mathrm{softmax}(W_{N}x)^{\top}\mathrm{softmax}(W_{N}y).$
5. Object‑class probes (pairwise) Alternatively, we can treat IsSameObject as a binary label and optimize $p^{\top}q$ directly with binary cross‑entropy loss. This version is pairwise because the supervision is applied directly to (x,y) as a pair. The model is thus optimized to encode the relationship between two patches, not their individual class memberships.
@@ 3.2 IsSameObject is best decodable in quadratic form
We extract DINOv2‑Large [13] activations at each layer and train the probes on the ADE20K dataset [56] using cross‑entropy loss for all pairwise probes to classify same‑object vs. different‑object patch pairs. Figure 2 shows probe accuracy across layers. To test our hypotheses about how IsSameObject is represented, we compare:
  • Linear vs. quadratic probes: Quadratic probes (even with the same parameter count as in diagonal quadratic) significantly outperform linear ones, suggesting that IsSameObject is a quadratic representation, consistent with the quadratic form used by the self-attention mechanism.
  • Quadratic vs. object-class probes: Trained object class probes fall short of the quadratic probe directly optimized for distinguishing same and different object instances. This indicates that the model encodes not just whether two patches share an object class but also subtler identity cues—e.g., distinguishing two identical cars of the same make and model.
  • Full vs. diagonal quadratic probes: The full quadratic probe outperforms its diagonal variant, implying that IsSameObject information is more distributed across dimensions rather than restricted to specific channels.
  • Pointwise object class probe vs. pairwise object class probe: Mapping each patch to an object class via softmax and then comparing classes (pointwise) underperforms directly comparing their embeddings (pairwise), as the pointwise approach loses information when it collapses embeddings into discrete classes.
@@ 3.3 Object binding emerges broadly across self-supervised ViTs
We extend our analysis beyond DINOv2 to a broader set of pretrained Vision Transformers, including CLIP, MAE, and fully supervised ViTs. To enable direct comparison, we standardize input patch coverage by resizing all inputs so that each model processes the same spatial patch divisions. Under this setup, every probe starts from the same trivial baseline of 72.6% accuracy, which corresponds to always predicting “different”, reflecting the class imbalance that most patch pairs do not belong to the same object in the dataset.
Table 1 reports IsSameObject decoding accuracy across models. DINO models show the strongest binding signal, with large and giant variants exceeding +16 percentage points over baseline. CLIP and MAE also exhibit clear object-binding ability, though to a lesser degree. In contrast, ImageNet-supervised ViT yields poor object‑binding performance, suggesting that binding is an acquired ability under specific pretraining objectives rather than being a universal property of all vision models.
Table 1: Binding is consistently represented in self-supervised ViTs, but less so in supervised ViTs. Probe accuracy on IsSameObject across pretrained ViTs. $\Delta$ is reported in percentage points ($p$), and the peak layer index is normalized to $[0,1]$ within each model.
@@
  • MAE. The masked autoencoder objective requires the model to reconstruct a missing patch from its surroundings. When the masked patch sits between multiple objects, correctly predicting its content forces the model to infer which object it belongs to, thus promoting the grouping of patches from the same object.
  • Supervised ImageNet training. Because ImageNet supervision labels only the dominant object in each image [57], it rewards class-level categorization but not grouping of patches into object instances, leading to a weaker binding signal.
@@ 4 Extracting the Binding Subspace of ViT Representations
Figure 3: The geometry of IsSameObject representation. Patch embeddings $h_{i}$ and $h_{j}$ are projected onto the IsSameObject subspace by $W_{probe}$, producing binding vectors $b_{i}$ and $b_{j}$, whose similarity is computed by a dot product.
@@ 4.1 Decomposing IsSameObject from features
Consider two identical patches $x_{A_{i}}$ and $x_{B_{i}}$ at corresponding positions of identical objects $A$ and $B$ in the same image, and let their residual be $\Delta_{AB_{i}}$. It may be tempting to cancel the feature term directly. Indeed, without positional encoding (see proof in Appendix A.4.1), we have $f^{(\ell)}(x_{A_{i}})=f^{(\ell)}(x_{B_{i}})$, because for identical tokens the positional encoding is the only signal that can differentiate their cross-token interactions.
@@
Instead, we take a supervised approach to decoding the binding component. Our quadratic probe serves as a tool for separating binding from feature information within each token (Fig. 3). Conceptually, the quadratic probe can be viewed as projecting an activation $h$ into the IsSameObject subspace, yielding $b^{(\ell)}(x)=h^{(\ell)}(x)^{\top}W$, and then measures the dot-product similarity between two projected vectors. Given that natural image datasets contain numerous objects where $b$ is the primary distinguishing factor, the probe should be optimized to discover a direction that isolates $b$. With this strategy we can separate the binding signal from the rest of the representation.
The observation that binding vectors remain meaningful under linear combination, and become hard to discriminate when they are close together, is consistent with this interpretation [52]. In later ablation studies, we use our trained quadratic probe via $b^{(\ell)}(x)=h^{(\ell)}(x)^{\top}W$.
@@ 4.2 A Toy Experiment: distinguishing identical objects and similar looking objects
To analyze where binding emerges, we plot the IsSameObject scores predicted by our trained quadratic probe (Figure 4). Since the probe performs reliably (see Section 3.2), its outputs approximate the internal IsSameObject the model encodes. We observe that, from early to mid‑layers, the model increasingly discerns the local object (the one to which each patch belongs). Surprisingly, from mid-layers to later layers, the model shifts toward class‑based grouping, increasingly treating all red cars as the same. Binding emerges in the middle of the network and is then progressively lost towards the top.
@@
We use four identical red‑car images and split each one into patches using exactly the same grid alignment. We perform principal component analysis (PCA) on the residuals sets $\{\Delta_{BA},\Delta_{CA},\Delta_{DA}\}$, where $\Delta_{BA}=h_{B_{i}}-h_{A_{i}}\approx b_{B_{i}}-b_{A_{i}}$ and visualize the first three components (see Figure 5). $\Delta_{BA},\Delta_{CA},\Delta_{DA}$ fall into three linearly separable clusters in the first three principal component space, showing that $b_{B},b_{C},b_{D}$ form object-level representations. The separation of these clusters in a very small number of principal directions demonstrates that IsSameObject lies in a low-dimensional subspace: patches from the same object instance map to closely aligned binding vectors, and different instances are linearly separable with large margins.
Figure 5: Identical objects form distinct object-level representations. The first 3 principal components of the four identical cars, with the three clusters denoting $\Delta_{BA},\Delta_{CA},\Delta{DA}$. The three linearly separable clusters suggest that identical objects’ binding vectors form distinct object‐level representations. The percentage in parentheses indicates the variance explained by that principal component.
@@
A surprising observation is the sudden increase in the cross-object IsSameObject score (fig.4) in the mid-layers of the DINOV2 model for instances of the same class (Fig. 4). This is consistent with prior work showing that ViTs represent different types of information at different layers [59]. At the same time, token-position decodability drops in deeper layers (see Appendix A.4.3), suggesting that the model is deliberately discarding positional information. Our interpretation is that the network initially relies on positional cues to support binding, since location is necessary to disambiguate tokens that share similar feature content. However, the DINO training objective enforces position invariance at the output layer [12], which implicitly encourages the network to remove positional signal once it is no longer useful and to repurpose capacity for semantically relevant object structure. Our findings are consistent with experimental evidence from the ventral stream in the brain, showing that while the retinotopic organization of early ventral areas is necessary for perception and binding, global spatial information is instead processed and maintained by the dorsal stream [60, 61, 62].

### v2

@@ Abstract
Object binding, the brain’s ability to bind the many features that collectively represent an object into a coherent whole, is central to human cognition. It groups low-level perceptual features into high‑level object representations, stores those objects efficiently and compositionally in memory, and supports human reasoning about individual object instances. While prior work often imposes object-centric attention (e.g., Slot Attention) explicitly to probe these benefits, it remains unclear whether this ability naturally emerges in pre-trained Vision Transformers (ViTs). Intuitively, they could: recognizing which patches belong to the same object should be useful for downstream prediction and thus guide attention. Motivated by the quadratic nature of self-attention, we hypothesize that ViTs represent whether two patches belong to the same object, a property we term IsSameObject. We decode IsSameObject from patch embeddings across ViT layers using a quadratic similarity probe, which reaches over 90% accuracy. Crucially, this object-binding capability emerges reliably in DINO, CLIP, and ImageNet-supervised ViTs, but is markedly weaker in MAE, suggesting that binding is not a trivial architectural artifact, but an ability acquired through specific pretraining objectives. We further discover that IsSameObject is encoded in a low-dimensional subspace on top of object features, and that this signal actively guides attention. Ablating IsSameObject from model activations degrades downstream performance and works against the learning objective, implying that emergent object binding naturally serves the pretraining objective. Our findings challenge the view that ViTs lack object binding and highlight how symbolic knowledge of “which parts belong together” emerges naturally in a connectionist system.  [^1] Code available at: https://github.com/liyihao0302/vit-object-binding.
@@ 1 Introduction
Object binding has received little attention in mainstream AI research. Cognition‑inspired models [27, 45] build in human-like object‑based attention. By contrast, mainstream vision models are assumed to implicitly learn to handle multiple objects from training data, yet empirical studies show they often "attend" only to the most salient regions and overlook the rest [24]. While ViT attention scores can capture global image structure and salient regions (often corresponding to target objects) [9], empirical evidence shows that self-attention tends to group patches by low-level feature similarity rather than reliably producing object-level binding [30]. Object‑centric methods like Slot Attention [29] fix this by allocating a small set of learnable slots that compete for token features, enforcing binding by design. However, whether AI vision models, especially leading ViTs, can achieve robust object binding without explicit mechanisms remains an open question.
Cognitive scientists have questioned whether ViTs can bind objects at all: arguing that they lack mechanisms for dynamically and flexibly grouping features [16]; they lack recurrence necessary for iterative refinement of object representations [35, 45]; and as purely connectionist models, they appear incapable of true symbolic processing [24]. However, these architectural limitations do not preclude binding from emerging through learning. If a model encodes whether two patches belong to the same object (IsSameObject), this signal can guide attention and improve prediction [9, 32]. Human-labeled data also reflects object-level structure, so ViTs can acquire binding by imitation. This suggests that ViTs may learn to bind objects directly from large-scale training data, without requiring explicit architectural inductive biases.
Here, we ask whether object binding naturally emerges in large, pretrained Vision Transformers, which is a question that matters for both cognitive science and AI. We propose IsSameObject (whether two patches belong to the same object) and show that it is reliably decodable (with 90.20% accuracy) using a quadratic similarity probe starting from mid-layers of the transformer layers. This effect is robust across DINO, CLIP and ImageNet-supervised ViTs, but largely absent in MAE, suggesting that binding is an acquired ability rather than a trivial architectural artifact. Across the ViT’s layer hierarchy, it progressively encodes IsSameObject in a low-dimensional projection-space on top of the features of the object, and it guides self-attention. Ablating IsSameObject from model activations hurts downstream performance and works against the pretraining objective.
@@ 2 Related Work
Motivated by how humans naturally reason about individual objects, Object-Centric Learning (OCL [29]) aims to represent a scene as a composition of disentangled object representations. While segmentation only partitions an image into object masks, OCL goes further by encoding each object into its own representation [15]. Unsupervised approaches such as MONet [7], IODINE [15], and especially Slot Attention [29] encode scenes into a small, permutation‑invariant set of “slots” that are iteratively refined, producing robust object representations on both synthetic [29, 26] and real‑world data [48, 50] and enabling compositional generation and manipulation [19, 20, 25]. However, since Slot Attention is added as an external module rather than integrated into the transformer architecture, it introduces additional challenges for scaling and training [42]. Other explicit object-centric approaches include Tensor Product Representations [51] and Capsule Networks [44].
Instead of object-centric approaches that explicitly enforce object-level attention, we propose an alternative view that ViTs may already encode implicit object‑level structure. Prior work has assumed this and attempted to group patches into objects directly from activations or attention maps ViTs, using methods like clustering [36] or GraphCut  [62].  [1] conduct a behavioral experiment where participants judge whether two dots belong to the same object at varying distances, and show that patch-level feature similarity in self-supervised ViTs supports object-based grouping. Building on this line of work, we show that ViT patch embeddings intrinsically encode whether any two patches belong to the same object, and analyze how this information is structured through probing.
@@ 3 Assessing Object Binding in ViTs through IsSameObject
Figure 2: IsSameObject predictions distinguish objects in highly complex scenarios [49]. Quadratic probe results for DINOv2-Large at layer 18 show that overlapping deer can be distinguished and that disconnected regions of the same deer are correctly retrieved in (c).
@@ 3.1 Probing IsSameObject representations
$\textbf{{IsSameObject}}\bigl(x_{i}^{(\ell)},x_{j}^{(\ell)}\bigr)=\phi\bigl(x_{i}^{(\ell)},x_{j}^{(\ell)}\bigr),\quad\phi:\mathbb{R}^{d}\times\mathbb{R}^{d}\to[0,1],$
@@
Here, we ask whether models reliably encode IsSameObject and, if so, what mechanisms they use to do so. We consider the following hypotheses about how IsSameObject may be encoded in the model’s activations:
@@
  • It is a pairwise relationship versus a pointwise mapping (i.e. the model first maps each patch to a discrete object identity or class, then compares).
  • The model tells objects apart using only broad class labels or object identities–i.e., it may rely on class-level recognition (“dog vs. chair”) instead of explicitly binding pixels to objects, as class labels already encode a coarse notion of object identity.
@@
To test these hypotheses, we decode IsSameObject using several probe architectures, each parameterized by a learnable matrix $W$ and a scalar bias $b$. All probes are constructed to be symmetric in their inputs, reflecting the constraint $\textit{IsSameObject}(x,y)=\textit{IsSameObject}(y,x)$. Throughout, $\sigma(\cdot)$ denotes the sigmoid function.
1. Linear probe.
$\textit{IsSameObject}_{\mathrm{lin}}(x,y)\;=\;\sigma\!\left(Wx+Wy+b\right),\qquad W\in\mathbb{R}^{1\times d},\;b\in\mathbb{R}.$
2. Diagonal quadratic probe (specialized dimensions).
$\textit{IsSameObject}_{\mathrm{diag}}(x,y)\;=\;\sigma\!\left(x^{\top}W\,y+b\right),$
where $W\in\mathbb{R}^{d\times d}$ is constrained to be diagonal, so the probe uses only $d$ parameters, each corresponding to a single feature dimension.
3. Quadratic probe (distributed).
$\textit{IsSameObject}_{\mathrm{quad}}(x,y)\;=\;\sigma\!\left(x^{\top}W_{1}^{\top}W_{2}\,y+b\right),$
where $W_{1},W_{2}\in\mathbb{R}^{k\times d}$ with $k\ll d$. To enforce symmetry, we set $W_{2}=SW_{1}$, where $S$ is a diagonal matrix with entries in $\{\pm 1\}$, yielding a low-rank quadratic form with $O(kd)$ parameters.
4. Object-class / object-identity probes (pointwise). We first map each embedding to a probability distribution:
$p=\mathrm{softmax}(W_{c}x+b),\qquad q=\mathrm{softmax}(W_{c}y+b),$
where $W_{c}$ is trained using multiclass cross-entropy on object-class labels (and similarly $W_{N}$ for object-identity labels). The pointwise IsSameObject score is then defined as the inner product of the two distributions:
$\textit{IsSameObject}_{\mathrm{class/identity}}(x,y)\;=\;p^{\top}q\;=\;\sum_{c=1}^{N_{c}}p(c)\,q(c).$
@@ 3.2 IsSameObject is best decodable in quadratic form
We extract DINOv2‑Large [32] activations at each layer and train the probes on the ADE20K dataset [64] using cross‑entropy loss for all pairwise probes to classify same‑object vs. different‑object patch pairs (see Figure 2 for IsSameObject visualizations). Figure 3 shows probe accuracy across layers. To test our hypotheses about how IsSameObject is represented, we compare:
  • Linear < quadratic probes: (Diagonal) quadratic probes significantly outperform linear ones, suggesting that IsSameObject is a quadratic representation, consistent with the quadratic form used by the self-attention mechanism.
  • Quadratic (pairwise) > object-identity probes (pointwise): Mapping each patch to a discrete object identity and then comparing them pointwise underperforms direct pairwise comparison of embeddings, as the pointwise approach discards information by collapsing continuous representations into discrete classes.
  • Quadratic > object class probes: The model encodes not only shared object class but also finer-grained identity cues (e.g., distinguishing two identical cars of the same make and model).
  • Full > diagonal quadratic probes: TheIsSameObject information is more distributed across dimensions rather than restricted to specific channels.
@@ 3.3 Object binding emerges broadly across self-supervised ViTs
We extend our analysis beyond DINOv2 to a broader set of pretrained Vision Transformers, including CLIP, MAE, and fully supervised ViTs. To enable direct comparison, we standardize input patch coverage by resizing all inputs so that each model processes the same spatial patch divisions as the DINOv2 family. Under this setup, every probe starts from the same trivial baseline of 72.6% accuracy, which corresponds to always predicting “different”, reflecting the class imbalance that most patch pairs do not belong to the same object in the dataset.
Table 1 reports IsSameObject decoding accuracy across models. DINO models show the strongest binding signal, with large and giant variants exceeding +16 percentage points over baseline. ImageNet-supervised ViT and CLIP also exhibit clear object-binding ability, though to a lesser degree. In contrast, MAE yields poor object‑binding performance, suggesting that binding is an acquired ability under specific pretraining objectives rather than being a universal property of all vision models.
Table 1: Binding is consistently represented in DINOv2, CLIP and supervised ViT, but less so in MAE. Probe accuracy on IsSameObject across pretrained ViTs. $\Delta$ is reported in percentage points ($p$), and the peak layer index is normalized to $[0,1]$ within each model.
@@
  • Supervised ImageNet training. Although ImageNet labels correspond to the dominant object in each image [43], class-level supervision still provides useful signals for object identity, consistent with the strong performance of our object-class probes.
@@ 4 Extracting the Binding Subspace of ViT Representations
Figure 4: The geometry of IsSameObject representation. Patch embeddings $h_{i}$ and $h_{j}$ are projected onto the IsSameObject subspace by $W_{1}$ (query) and $W_{2}$ (key), producing binding vectors, whose similarity is computed by a dot product. The asymmetry between query and key spaces (where $W_{2}$ differs from $W_{1}$ by sign-flipped rows) mirrors the asymmetric roles of query and key in self-attention.
@@ 4.1 Decomposing IsSameObject from features
Consider two identical patches $x_{A_{i}}$ and $x_{B_{i}}$ at corresponding positions of identical objects $A$ and $B$ in the same image, and let their residual be $\Delta_{AB_{i}}$. It may be tempting to cancel the feature term directly. Indeed, without positional encoding (see proof in Appendix A.4.1), we have $f^{(\ell)}(x_{A_{i}})=f^{(\ell)}(x_{B_{i}})$, since for identical tokens the positional encoding is the only signal that can differentiate their cross-token interactions.
@@
Instead, we take a supervised approach to decoding the binding component. Our quadratic probe serves as a tool for separating binding from feature information within each token (Fig. 4). Conceptually, the quadratic probe can be viewed as projecting an activation $h$ into the IsSameObject subspace, yielding $b_{query}^{(\ell)}(x)=h^{(\ell)}(x)^{\top}W_{1}$ and $b_{key}^{(\ell)}(x)=h^{(\ell)}(x)^{\top}W_{2}$, and then measures the dot-product similarity between two projected vectors. Given that natural image datasets contain numerous objects where $b$ is the primary distinguishing factor, the probe should be optimized to discover a direction that isolates $b$. With this strategy we can separate the binding signal from the rest of the representation.
The observation in [14] that binding vectors remain meaningful under linear combination, and become hard to discriminate when they are close together, is consistent with this interpretation. In later ablation studies, we use our trained quadratic probe via $b^{(\ell)}(x)=h^{(\ell)}(x)^{\top}W$.
@@ 4.2 A Toy Experiment: distinguishing identical objects and similar looking objects
To analyze where binding emerges, we plot the IsSameObject scores predicted by our trained quadratic probe (Figure 5). We observe that, from early to mid‑layers, the model increasingly discerns the local object (the one to which each patch belongs). Surprisingly, from mid-layers to later layers, the model shifts toward class‑based grouping, increasingly treating all red cars as the same. Binding emerges in the middle of the network and is then progressively lost towards the top.
@@
We use four identical red‑car images and split each one into patches using exactly the same grid alignment. We perform principal component analysis (PCA) on the residuals sets $\{\Delta_{BA},\Delta_{CA},\Delta_{DA}\}$, where $\Delta_{BA}=h_{B_{i}}-h_{A_{i}}\approx b_{B_{i}}-b_{A_{i}}$ and visualize the first three components (see Figure 6). $\Delta_{BA},\Delta_{CA},\Delta_{DA}$ fall into three linearly separable clusters in the first three principal component space. The separation of these clusters in a very small number of principal directions demonstrates that IsSameObject lies in a low-dimensional subspace: patches from the same object instance map to closely aligned binding vectors, and different instances are linearly separable with large margins.
Figure 6: Identical objects form distinct object-level representations. The first 3 principal components of the four identical cars, with the three linearly separable clusters denoting $\Delta_{BA},\Delta_{CA},\Delta{DA}$. The percentage in parentheses indicates the variance explained by that principal component.
@@
A surprising observation is the sudden increase in the cross-object IsSameObject score (Fig.5) in the mid-layers of the DINOV2 model for instances of the same class (Fig. 5). This is consistent with prior work showing that ViTs represent different types of information at different layers [3]. At the same time, token-position decodability drops in deeper layers (see Appendix A.4.3), suggesting that the model is deliberately discarding positional information. Our interpretation is that the network initially relies on positional cues to support binding, since location is necessary to disambiguate tokens that share similar feature content. In later layers, the network removes positional signals once they are no longer useful and repurposes capacity for semantically relevant object structure. Our findings are consistent with experimental evidence from the ventral stream in the brain, showing that while the retinotopic organization of early ventral areas is necessary for perception and binding, global spatial information is instead processed and maintained by the dorsal stream [4, 56, 5].